OpenStudy (anonymous):

lim x -->1 cuberoot of x-1/ sqrt of x-1

OpenStudy (anonymous):

well what do you think the answer is

OpenStudy (anonymous):

i got 2...but I am not sure

OpenStudy (solomonzelman):

\(\color{#000000}{\displaystyle\lim_{x \rightarrow ~1} \frac{\sqrt[3]{x-1}}{\sqrt{x-1}} }\) like this?

OpenStudy (anonymous):

yes :-)

OpenStudy (anonymous):

no no no

OpenStudy (anonymous):

it's 2

OpenStudy (anonymous):

the cube root and square root cover only x

OpenStudy (solomonzelman):

\(\color{#000000}{\displaystyle\lim_{x \rightarrow ~1} \frac{\sqrt[3]{x}-1}{\sqrt{x}-1} }\)

OpenStudy (solomonzelman):

this?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

Differentiate top and bottom, by L'H'S

OpenStudy (solomonzelman):

So far, if you plug in x=0 into the limit, you get 0/0, therefore, you may apply L'Hospital's rule in this case,

OpenStudy (anonymous):

I am not allowed to use l'hopital's rule

OpenStudy (solomonzelman):

Interesting ...

OpenStudy (anonymous):

yeah...I tried rationalizing but my denominator is still zero no matter what

OpenStudy (anonymous):

I got 2 on my last try but I am still not sure 'cause i took a very sketchy approach

OpenStudy (solomonzelman):

I did some rationalization, and I am getting... \(\color{#000000 }{ \displaystyle \lim_{x\to 1}\frac{x^{5/6}-x^{1/2}+x^{1/3}-1}{x-1} }\) not ure how that's useful

myininaya (myininaya):

try a sub let x=u^6

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \lim_{x\to 1}\frac{x^{1/3}-1}{(x^{1/3})^3-1} }\) \(\color{#000000 }{ \displaystyle \lim_{x\to 1}\frac{x^{1/3}-1}{(x^{1/3}-1)(x^{2/3}+x^{1/3}+1)} }\)

OpenStudy (solomonzelman):

Using, \(\color{#000000 }{ \displaystyle a^3-b^3=(a-b)(a^2+ab+b^2) }\)

OpenStudy (solomonzelman):

then cancel \(\color{#000000 }{ \displaystyle (x^{1/3}-1) }\)

OpenStudy (anonymous):

I thought it should be x-1-1 on the numerator after rationalizing

myininaya (myininaya):

you rewrote the denominator as x-1 @SolomonZelman ?

myininaya (myininaya):

Also if you make the sub x=u^6 your problem becomes much easier to deal with

OpenStudy (solomonzelman):

I wrote, \(\color{#000000 }{ \displaystyle x-1=(x^{1/3})^3-1}\)

OpenStudy (solomonzelman):

Then, used the rule \(a^3-b^3\) ... blah blah blah

OpenStudy (anonymous):

I am so confused!

OpenStudy (solomonzelman):

What are you confused about?

myininaya (myininaya):

but sqrt(x)-1 isn't x-1

OpenStudy (anonymous):

I used that rule and after rationalizing I got...

OpenStudy (solomonzelman):

oh, I see... my bad

OpenStudy (solomonzelman):

I did it as if it is that, when it isn't that, but this .. anyway:)

myininaya (myininaya):

\[\lim_{u \rightarrow 1}\frac{u^2-1}{u^3-1}\] after the sub I suggested factor a bit and cancel

OpenStudy (solomonzelman):

Yeah, that works:)

OpenStudy (anonymous):

\[((\sqrt[3]{x}-1)\div (\sqrt{x}-1)) \times ((\sqrt[3]{x})^{2}+\sqrt[3]{x}+1\div(\sqrt[3]{x})^{2}+\sqrt[3]{x}+1)\]

OpenStudy (anonymous):

that's the rationalization right???

OpenStudy (solomonzelman):

why did you divide by the denominator?

OpenStudy (anonymous):

no...I applied...(a^3)-b^3....rule

OpenStudy (anonymous):

The original expression is on one side...and the 'conjugate' is on the right side

myininaya (myininaya):

if you are so fixed on a rationalization method over the sub method... then try... \[\frac{x^\frac{1}{3}-1}{x^\frac{1}{2}-1} \cdot \frac{x^\frac{1}{2}+1}{x^\frac{1}{2}+1} \cdot \frac{x^\frac{2}{3}+x^{\frac{1}{3}}+1}{x^\frac{2}{3}+x^\frac{1}{3}+1} \\ =\frac{(x^\frac{1}{3}-1)(x^\frac{2}{3}+x^\frac{1}{3}+1)}{(x^\frac{1}{2}-1)(x^\frac{1}{2}+1)} \cdot \frac{x^\frac{1}{2}+1}{x^\frac{2}{3}+x^\frac{1}{3}+1}\]

myininaya (myininaya):

it is like sorta rationalizing top and bottom (while really not; you are just doing it to cancel a factor )

myininaya (myininaya):

that first fraction is a rational you will see this after preforming the multiplication on top and bottom

myininaya (myininaya):

honestly the sub mentioned earlier makes finding the limit easier than trying to rationalize (my opinion)

OpenStudy (anonymous):

Thank you guys so much for even replying in the first place...I really appreciate your help :-) @myininaya @SolomonZelman

OpenStudy (solomonzelman):

Yes, indeed. \(\color{#000000 }{ \displaystyle \lim_{x\to 1}\frac{x^{1/n}-1}{x^{1/2}-1} }\) Will make a substitution: \(\color{#000000 }{ \displaystyle x=z^{2n} }\) Note that as \(\color{#000000 }{ \displaystyle x \to 1 }\), \(\color{#000000 }{ \displaystyle z \to 1 }\). \(\color{#000000 }{ \displaystyle \lim_{z\to 1}\frac{(z^{2n})^{1/n}-1}{(z^{2n})^{1/2}-1}= \lim_{z\to 1}\frac{z^2-1}{z^n-1}=\lim_{z\to 1}\frac{2z}{nz^{n-1}} =\frac{2}{n} }\)

OpenStudy (solomonzelman):

That is the general derivation through L'H'S, not that it is applicable in your case, because it is not premitted as you said.

OpenStudy (anonymous):

okay...I think i will just try the substitution method..

OpenStudy (anonymous):

Thank you!!!:-)

OpenStudy (solomonzelman):

In your example (\(x=u^6\), as myininaya has offered). In that case, the top and bottom are factor-able ...

OpenStudy (solomonzelman):

GO for it... yw

OpenStudy (anonymous):

Okay...Thanks a bunch! @myininaya @SolomonZelman

OpenStudy (solomonzelman):

Not a problem!

OpenStudy (anonymous):

:-)